# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v2

@@ 4 DISTRIBUTIONALLY AND ADVERSARIALLY ROBUST LR
By employing a simple duality trick for the inner $\sup$-problem, as commonly applied in robust optimization (Ben-Tal et al.,, 2009; Bertsimas and Den Hertog,, 2022), we can represent DR-ARO as a standard non-adversarial DRO problem with an updated loss function, which we name the adversarial loss.
@@
As a corollary of Observation 1, we can directly employ the techniques proposed by Shafieezadeh-Abadeh et al., (2019) to dualize the inner $\sup$-problem of DR-ARO and obtain a tractable reformulation.
@@
$\begin{array}[]{cll}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\bm{s}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\dfrac{1}{N}\sum_{i=1}^{N}s_{i}\\[11.99998pt] \mathrm{s.t.}&\displaystyle\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},y^{i})\leq s_{i}&\forall i\in[N]\\[1.99997pt] &\displaystyle\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},-y^{i})-\lambda\kappa\leq s_{i}&\forall i\in[N]\\[1.99997pt] &\displaystyle\lVert\bm{\beta}\rVert_{q^{\star}}\leq\lambda\\[1.99997pt] &\displaystyle\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\end{array}$
@@
The constraints of this problem are exponential cone representable (derivation is in the appendices) and for $q\in\{1,2,\infty\}$, the yielding problem can be solved with the exponential cone solver of MOSEK (MOSEK ApS,, 2023) in polynomial time (Nesterov,, 2018).
@@ 5 MAIN RESULT
The assumption implies that we know $\varepsilon,\widehat{\varepsilon}>0$ values satisfying $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$. In practice, this assumption is ensured to hold with high confidence by estimating the $\varepsilon$ and $\widehat{\varepsilon}$ values; methods across various domains which we can adopt are reviewed in §6. Under Assumption 2, we want to optimize the adversarial loss over the intersection $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$:
@@
This formulation is expected to outperform DR-ARO as the ambiguity set is smaller while still including $\mathbb{P}^{0}$. However, problem Inter-ARO is challenging to solve even in the absence of adversarial attacks ($\alpha=0$) as we reviewed in §2. To address this challenge, we first reformulate Inter-ARO as a semi-infinite optimization problem with finitely many variables.
@@
$\begin{array}[]{cl}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\widehat{\lambda}\\ \bm{s},\widehat{\bm{s}}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\left[\begin{aligned} \underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},l)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\\ \leq s_{i}+\frac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\frac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda}\end{aligned}\right]\\[14.22636pt] &\hfill\forall i\in[N],\ j\in[\widehat{N}],\ l\in\{-1,1\}\\[2.84526pt] &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}.\end{array}$
Even though this problem recovers the tractable problem DR-ARO as $\widehat{\varepsilon}\rightarrow\infty$, it is NP-hard in the finite radius settings. We reformulate Inter-ARO as an adjustable robust optimization problem (Ben-Tal et al.,, 2004; Yanıkoğlu et al.,, 2019), and borrow tools from this literature to obtain the following result.
@@
The adjustable RO literature has developed a rich arsenal of relaxations that can be leveraged for Inter-ARO. We adopt the ‘static relaxation technique’ (Bertsimas et al.,, 2015) to restrict the feasible region of Inter-ARO and obtain a tractable approximation.
@@
$\begin{array}[]{cl}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}}\\ \bm{z}_{ij}^{+},\bm{z}_{ij}^{-}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\left[\begin{aligned} &L^{\alpha}(l\cdot\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}_{ij}^{l\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))\\ &\ \leq s_{i}+\dfrac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\dfrac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda},\\[3.41432pt] &\lVert l\bm{\beta}-\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\lambda,\;\lVert\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\widehat{\lambda}\end{aligned}\right]\\[22.76219pt] &\hfill\forall i\in[N],\ j\in[\widehat{N}],\ l\in\{-1,1\}\\ &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}\\ &\bm{z}^{l}_{ij}\in\mathbb{R}^{n},\ (i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}.\end{array}$ (16)
Similarly to DR-ARO, the constraints of 16 are exponential cone representable (cf. appendices).
Recall that for $\widehat{\varepsilon}$ large enough, Inter-ARO reduces to DR-ARO. The following corollary shows that, despite 16 being a relaxation of Inter-ARO, a similar property holds. That is, “not learning anything from auxiliary data” remains feasible: the static relaxation does not force learning from ^$\widehat{\mathbb{P}}_{\widehat{N}}$, and it learns from auxiliary data only if the objective improves.
@@
Feasibility of ignoring auxiliary data: Any feasible solution $(\bm{\beta},\lambda,\bm{s})$ of DR-ARO can be used to recover a feasible solution $(\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}},\bm{z}^{+}_{ij},\bm{z}^{-}_{ij})$ for 16 with $\widehat{\lambda}=0$, $\widehat{\bm{s}}=\mathbf{0}$, and $\bm{z}^{+}_{ij}=\bm{z}^{-}_{ij}=\mathbf{0}$.
Convergence to Inter-ARO: The optimal value of 16 converges to the optimal value of Inter-ARO, with the same set of $\bm{\beta}$ solutions, as $\widehat{\varepsilon}\rightarrow\infty$.
We close the section by discussing how Inter-ARO can recover some problems in the DRO and ARO literature. Firstly, recall that Inter-ARO can ignore the auxiliary data once $\widehat{\varepsilon}$ is set large enough, reducing this problem to DR-ARO. Moreover, notice that $\alpha=0$ reduces $\ell_{\bm{\beta}}^{\alpha}$ to $\ell_{\bm{\beta}}$, hence for $\alpha=0$ and $\widehat{\varepsilon}=\infty$ Inter-ARO recovers the Wasserstein LR model of Shafieezadeh-Abadeh et al., (2015). We next relate Inter-ARO to the problems in the ARO literature that use auxiliary data. The works in this literature (Gowal et al.,, 2021; Xing et al., 2022b, ) solve the following for some $w>0$:
@@
where $\mathcal{B}_{p}(\alpha):=\{\bm{z}\in\mathbb{R}^{n}:\lVert\bm{z}\rVert_{p}\leq\alpha\}$. We first observe that this resembles 8, with the empirical distribution $\mathbb{P}_{N}$ being replaced with its mixture with ^$\widehat{\mathbb{P}}_{\widehat{N}}$:
@@
For $\lambda=\frac{N}{N+\widehat{N}}$, if the intersection $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ is nonempty, Proposition 3 implies that a sufficient condition for this intersection to include $\mathbb{Q}_{\mathrm{mix}}$ is $\widehat{\varepsilon}/\varepsilon=N/\widehat{N}$, which is intuitive since the radii of Wasserstein ambiguity sets are chosen inversely proportional to the number of samples (Kuhn et al.,, 2019, Theorem 18).
@@ 6 SETTING WASSERSTEIN RADII
Theorem 2 (abridged).
@@
Choosing $\epsilon$ and $\epsilon^{\prime}$ in Inter-ARO. Recall that Inter-ARO revises DR-ARO by intersecting $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with $\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$. We need a nonempty intersection for Inter-ARO to be well-defined. A sufficient condition follows from the triangle inequality: $\varepsilon+\widehat{\varepsilon}\geq\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$. We also want this intersection to include $\mathbb{P}^{0}$ with high confidence, in order to satisfy Assumption 2. We next provide a tight characterization for such $\varepsilon,\widehat{\varepsilon}$. The full description of this result is available in our appendices.
@@
In Theorem 3, we use $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ explicitly. This distance, however, is typically unknown, and a common approach is to cross-validate it [^1] Note that, in practice, the distance between the unknown true and auxiliary data-generating distributions is also cross-validated in the transfer learning and domain adaptation literature (Zhong et al.,, 2010).. This would be applicable in our setting thanks to Corollary 2, because the relaxation problem 16 is not forced to learn from the auxiliary data unless it is useful, that is, one can find evidence for the usefulness of the auxiliary data via cross-validation. Moreover, there are several domains where $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ is known exactly. For some special cases, we can use direct domain knowledge (e.g., the “Uber vs Lyft” example of Taskesen et al., 2021). A very recent example comes from learning from multi-source data, where $\mathbb{P}^{0}$ is named the target distribution and $\widehat{\mathbb{P}}$ is the source distribution (Rychener et al.,, 2024, §1). Another domain is private data release, where a data holder shares a subset of opt-in data to form $\mathbb{P}_{N}$, and generates a privacy-preserving synthetic dataset from the rest. The (privately generated) synthetic distribution has a known nonzero Wasserstein distance from the true data-generating distribution (Dwork et al.,, 2014; Ullman and Vadhan,, 2020). Alternatively, one can directly rely on $\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}})$ when it is known, especially when synthetic data generators are trained on the empirical dataset. By employing Wasserstein GANs, which minimize the Wasserstein-1 distance, the distance between the generated distribution and the training distribution is minimized. This ensures that the synthetic distribution remains within a small radius of the training distribution (Arjovsky et al.,, 2017).

### v4

@@ 4 DISTRIBUTIONALLY AND ADVERSARIALLY ROBUST LR
By employing a simple duality trick for the inner $\sup$-problem, as commonly applied in robust optimization (Ben-Tal et al., 2009; Bertsimas and Den Hertog, 2022), we can represent DR-ARO as a standard non-adversarial DRO problem with an updated loss function, which we name the adversarial loss.
@@
As a corollary of Observation 1, we can directly employ the techniques proposed by Shafieezadeh-Abadeh et al. (2019) to dualize the inner $\sup$-problem of DR-ARO and obtain a tractable reformulation.
@@
$\begin{array}[]{cll}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\bm{s}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\dfrac{1}{N}\sum_{i=1}^{N}s_{i}\\[11.99998pt] \mathrm{s.t.}&\displaystyle\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},y^{i})\leq s_{i}&\forall i\in[N]\\[6.00006pt] &\displaystyle\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},-y^{i})-\lambda\kappa\leq s_{i}&\forall i\in[N]\\[6.00006pt] &\displaystyle\lVert\bm{\beta}\rVert_{q^{\star}}\leq\lambda\\[6.00006pt] &\displaystyle\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\end{array}$
@@
The constraints of this problem are exponential cone representable (derivation is in the appendices) and for $q\in\{1,2,\infty\}$, the yielding problem can be solved with the exponential cone solver of MOSEK (MOSEK ApS, 2023) in polynomial time (Nesterov, 2018).
@@ 5 MAIN RESULT
The assumption implies that we know $\varepsilon,\widehat{\varepsilon}>0$ values satisfying $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$. In practice, this assumption is ensured to hold with high confidence by estimating the $\varepsilon$ and $\widehat{\varepsilon}$ values; methods across various domains which we can adopt are reviewed in §6. We want to optimize the adversarial loss over the intersection $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$:
@@
Note that Assumption 2 implies that $\varepsilon$ and $\widehat{\varepsilon}$ guarantee that $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$ is nonempty, and problem Inter-ARO is thus feasible. This problem is expected to outperform DR-ARO as the ambiguity set is smaller while still including $\mathbb{P}^{0}$. However, problem Inter-ARO is challenging to solve even in the absence of adversarial attacks ($\alpha=0$) as we reviewed in §2. To address this challenge, we first reformulate Inter-ARO as a semi-infinite optimization problem with finitely many variables.
@@
$\begin{array}[]{cl}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\widehat{\lambda}\\ \bm{s},\widehat{\bm{s}}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\dfrac{1}{N}\sum_{i=1}^{N}s_{i}+\dfrac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},l)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\\ &\hfill\leq s_{i}+\dfrac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\dfrac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda}\\ &\hfill\forall(i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}\\[2.84526pt] &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}.\end{array}$
Even though this problem recovers the tractable problem DR-ARO as $\widehat{\varepsilon}\rightarrow\infty$, it is NP-hard in the finite radius settings. We reformulate Inter-ARO as an adjustable robust optimization problem (Ben-Tal et al., 2004; Yanıkoğlu et al., 2019), and borrow tools from this literature to obtain the following result.
@@
The adjustable RO literature has developed a rich arsenal of relaxations that can be leveraged for Inter-ARO. We adopt the ‘static relaxation technique’ (Bertsimas et al., 2015) to restrict the feasible region of Inter-ARO and obtain a tractable approximation.
@@
$\begin{array}[]{cl}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}}\\ \bm{z}_{ij}^{+},\bm{z}_{ij}^{-}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&L^{\alpha}(l\cdot\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}_{ij}^{l\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))\\ &\hfill\leq s_{i}+\dfrac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\dfrac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda}\\ &\hfill\forall(i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}\\[8.00003pt] &\lVert l\bm{\beta}-\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\lambda\\ &\hfill\forall(i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}\\[8.00003pt] &\lVert\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\widehat{\lambda}\\ &\hfill\forall(i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}\\[8.00003pt] &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}\\[8.00003pt] &\bm{z}^{l}_{ij}\in\mathbb{R}^{n},\ (i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}.\end{array}$ (21)
Similarly to DR-ARO, the constraints of 21 are exponential cone representable (cf. appendices).
Recall that for $\widehat{\varepsilon}$ large enough, Inter-ARO reduces to DR-ARO. The following corollary shows that, despite 21 being a relaxation of Inter-ARO, a similar property holds. That is, “not learning anything from auxiliary data” remains feasible: the static relaxation does not force learning from ^$\widehat{\mathbb{P}}_{\widehat{N}}$, and it learns from auxiliary data only if the objective improves.
@@
Feasibility of ignoring auxiliary data: Any feasible solution $(\bm{\beta},\lambda,\bm{s})$ of DR-ARO can be used to recover a feasible solution $(\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}},\bm{z}^{+}_{ij},\bm{z}^{-}_{ij})$ for 21 with $\widehat{\lambda}=0$, $\widehat{\bm{s}}=\mathbf{0}$, and $\bm{z}^{+}_{ij}=\bm{z}^{-}_{ij}=\mathbf{0}$.
Convergence to Inter-ARO: The optimal value of 21 converges to the optimal value of Inter-ARO, with the same set of $\bm{\beta}$ solutions, as $\widehat{\varepsilon}\rightarrow\infty$.
In light of Corollary 2, Appendix D.1 discusses that some simulations in our numerical experiments chose not to incorporate the auxiliary data by setting a sufficiently large $\widehat{\varepsilon}$. We close the section by discussing how Inter-ARO can recover some problems in the DRO and ARO literature. Firstly, recall that Inter-ARO can ignore the auxiliary data once $\widehat{\varepsilon}$ is set large enough, reducing this problem to DR-ARO. Moreover, notice that $\alpha=0$ reduces $\ell_{\bm{\beta}}^{\alpha}$ to $\ell_{\bm{\beta}}$, hence for $\alpha=0$ and $\widehat{\varepsilon}=\infty$ Inter-ARO recovers the Wasserstein LR model of Shafieezadeh-Abadeh et al. (2015). We next relate Inter-ARO to the problems in the ARO literature that use auxiliary data. The works in this literature (Gowal et al., 2021; Xing et al., 2022a) solve the following
@@
for some $w>0$, where $\mathcal{B}_{p}(\alpha):=\{\bm{z}\in\mathbb{R}^{n}:\lVert\bm{z}\rVert_{p}\leq\alpha\}$. We observe that (24) resembles a variant of 8 that replaces the empirical distribution $\mathbb{P}_{N}$ with its mixture with ^$\widehat{\mathbb{P}}_{\widehat{N}}$:
@@
For $\lambda=\frac{N}{N+\widehat{N}}$, if the intersection $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ is nonempty, Proposition 3 implies that a sufficient condition for this intersection to include $\mathbb{Q}_{\mathrm{mix}}$ is $\widehat{\varepsilon}/\varepsilon=N/\widehat{N}$, which is intuitive since the radii of Wasserstein ambiguity sets are chosen inversely proportional to the number of samples (Kuhn et al., 2019, Theorem 18).
@@ 6 SETTING WASSERSTEIN RADII
Theorem 2 (abridged collection of results from Fournier and Guillin 2015; Kuhn et al. 2019; Yue et al. 2022).
@@
Choosing $\epsilon$ and $\widehat{\varepsilon}$ in Inter-ARO. Recall that Inter-ARO revises DR-ARO by intersecting $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with $\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$. We need a nonempty intersection for Inter-ARO to be well-defined. A necessary and sufficient condition follows from the triangle inequality $\varepsilon+\widehat{\varepsilon}\geq\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$, where $\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$ can be computed with linear optimization as both distributions are discrete. We also want this intersection to include $\mathbb{P}^{0}$ with high confidence, in order to satisfy Assumption 2. We next provide a tight characterization for such $\varepsilon,\widehat{\varepsilon}$. The full description of this result is available in our appendices.
@@
Remark 2.
The assumption that true data-generating distributions are light-tailed is satisfied when $\Xi$ is compact, and it is a common assumption for even simple sample average approximation techniques (Mohajerin Esfahani and Kuhn, 2018, Assumption 3.3).
@@
In Theorem 3, we use $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ explicitly. This distance, however, is typically unknown, and a common approach is to cross-validate it [^1] In practice, distance between the unknown true and auxiliary data-generating distributions is also cross-validated in the transfer learning and domain adaptation literature (Zhong et al., 2010).. This would be applicable in our setting thanks to Corollary 2, because the relaxation 21 does not force learning from the auxiliary data unless it is useful, that is, one can seek evidence for the usefulness of the auxiliary data via cross-validation. Moreover, there are several domains where $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ is known exactly. For some special cases, we can use direct domain knowledge (e.g., the “Uber vs Lyft” example of Taskesen et al. 2021). A recent example comes from learning from multi-source data, where $\mathbb{P}^{0}$ is named the target distribution and $\widehat{\mathbb{P}}$ is the source distribution (Rychener et al., 2024, §1). Another domain is private data release, where a data holder shares a subset of opt-in data to form $\mathbb{P}_{N}$, and generates a privacy-preserving synthetic dataset from the rest. The (privately generated) synthetic distribution has a known nonzero Wasserstein distance from the true data-generating distribution (Dwork and Roth, 2014; Ullman and Vadhan, 2020). See (Rychener et al., 2024, §5) for more scenarios that enable quantifying $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$. Alternatively, one can directly rely on $\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}})$ if it is known, especially when synthetic data generators are trained on the empirical dataset. By employing Wasserstein GANs, which minimize the Wasserstein-1 distance, the distance between the generated distribution and the training distribution is minimized. This ensures that the synthetic distribution remains within a radius of the training distribution (Arjovsky et al., 2017).